Melinda invested three times as much money at 9% yearly interest as she did at 7%. Her total yearly interest from the two investments was $170. How much did she invest at each rate?
I only set up this
9(.09x)+ .07x= 170

Answer by stanbon(75887)

Melinda invested three times as much money at 9% yearly interest as she did at 7%. Her total yearly interest from the two investments was $170. How much did she invest at each rate?
------
Quantity Equation: n = 3s
---
Interest Eqution:: 0.09n + 0.07s = 170
---------
Substitute for "n" and solve for "s"
0.09(3s) + 0.07s = 170
Multiply thru by 100 to get:
27s + 7s = 17000
34s = 17000
s = $500.00 (amt. invested at 7%)
----
solve for "n" using n = 3s
n = $1500 (amt. invested at 9%)
===================
Cheers,
stan H.
================